# Dynamics of the general $Q$-tensor model interacting with a rigid body

### v1

@@ 7. Local strong well-posedness for large data
$\|\Phi_{T}^{R}(\tilde{z}_{1},\tilde{p}_{1})-\Phi_{T}^{R}(\tilde{z}_{2},\tilde{p}_{2})\|_{\tilde{\mathbb{E}}_{T}}\leq C_{\mathrm{MR}}L(R,T)\cdot\|(\tilde{z}_{1},\tilde{p}_{1})-(\tilde{z}_{2},\tilde{p}_{2})\|_{\tilde{\mathbb{E}}_{T}},\enspace\text{for}\enspace(\tilde{z}_{1},\tilde{p}_{1}),\,(\tilde{z}_{2},\tilde{p}_{2})\in\mathcal{K}_{T}^{R}.$
As $C_{\mathrm{MR}}>0$ is $T$-independent by virtue of the homogeneous initial values, we may choose $R>0$ and then also $T>0$ sufficiently small such that $C_{\mathrm{MR}}C(R,T)\leq\nicefrac{{R}}{{2}}$ and $C_{\mathrm{MR}}L(R,T)\leq\nicefrac{{1}}{{2}}$. In other words, the solution map $\Phi_{T}^{R}$ to the linearized problem (7.1) is a self-map and contraction, giving rise to a unique fixed point $(\widehat{z},\widehat{p})\in\prescript{}{0}{\mathbb{E}_{T}}\times\mathbb{E}_{T}^{p}$. An addition of the reference solution $(z^{*},p^{*})$ from Section 7 leads to the unique solution $(z,p)=(\widehat{z}+z^{*},\widehat{p}+p^{*})\in\tilde{\mathbb{E}}_{T}$ to (5.11)–(5.13). The proof is completed upon noting that $(z,p)\in\tilde{\mathbb{E}}_{T}$ means that $(v,Q,\ell,\omega,p)$ is in the regularity class as stated in Theorem 5.1.
@@ 8. Global strong well-posedness for small data
(8.1) $\partial_{t}{v}-\Delta{v}+\nabla{p}-\operatorname{div}S_{\xi}(0)(\Delta-\operatorname{Id}_{3})Q&=G_{1}(z,p),\enspace\operatorname{div}{v}=0,&&\enspace\text{in}\enspace(0,T)\times\mathcal{F}_{0},\\ {\partial_{t}{Q}}+\tilde{S}_{\xi}(0)\nabla v-(\Delta-\operatorname{Id}_{3}){Q}&=G_{2}(z),&&\enspace\text{in}\enspace(0,T)\times\mathcal{F}_{0},\\ m_{\mathcal{S}}({\ell})^{\prime}+\int_{\partial\mathcal{S}_{0}}\mathrm{T}(v,p,Q)n\,\mathrm{d}\Gamma&=G_{3}(z,p),&&\enspace\text{in}\enspace(0,T),\\ J_{0}({\omega})^{\prime}+\int_{\partial\mathcal{S}_{0}}y\times\mathrm{T}(v,p,Q)n\,\mathrm{d}\Gamma&=G_{4}(z,p),&&\enspace\text{in}\enspace(0,T),\\ {v}={\ell}+{\omega}\times y,\enspace\text{on}\enspace(0,T)\times\partial\mathcal{S}_{0},\enspace{v}&=0,\enspace\text{on}\enspace(0,T)\times\partial\mathcal{O},\enspace\partial_{\nu}Q=0,&&\enspace\text{on}\enspace(0,T)\times\partial\mathcal{F}_{0},\\ v(0)=v_{0},\enspace Q(0)=Q_{0},\enspace\ell(0)&=\ell_{0}\enspace\text{and}\enspace\omega(0)=\omega_{0}.\end{aligned}\right.$
@@
(8.2) $G_{1}(z,p)$ $\coloneqq(\mathcal{L}_{1}-\Delta)v-\mathcal{M}v-\mathcal{N}(v)-(\mathcal{G}-\nabla)p+\mathcal{B}_{\tau}(Q)+\mathcal{B}_{\sigma}(Q)-\operatorname{div}S_{\xi}(0)(\Delta-\operatorname{Id}_{3})Q,$
@@
$G_{3}(z,p)$ $\coloneqq-m_{\mathcal{S}}\omega\times\ell+\int_{\partial\mathcal{S}_{0}}\bigl{(}(\mathrm{T}-\mathcal{T})(v,p,Q)-\mathbb{O}(t)^{\top}(\tau_{\mathrm{r}}(Q)+\sigma(Q))\mathbb{O}(t)\bigr{)}n\,\mathrm{d}\Gamma\enspace\text{and}\enspace$
$G_{4}(z,p)$ $\coloneqq\omega\times(J_{0}\omega)+\int_{\partial\mathcal{S}_{0}}y\times\bigl{(}(\mathrm{T}-\mathcal{T})(v,p,Q)-\mathbb{O}(t)^{\top}(\tau_{\mathrm{r}}(Q)+\sigma(Q))\mathbb{O}(t)\bigr{)}n\,\mathrm{d}\Gamma.$
For $\eta\in(0,\eta_{0})$, where $\eta_{0}>0$ has been introduced in Theorem 6.1(b), and $\tilde{\mathbb{E}}_{\infty}$ as made precise in (6.3), we define $\tilde{\mathcal{K}}$ to be $\tilde{\mathbb{E}}_{\infty}$ with exponential weights, i. e., $(z,p)=(v,Q,\ell,\omega,p)\in\tilde{\mathcal{K}}$ if and only if
(8.3) $\mathrm{e}^{\eta(\cdot)}(v,Q,\ell,\omega,p)\in\tilde{\mathbb{E}}_{\infty},\enspace\text{with}\enspace\|(v,Q,\ell,\omega,p)\|_{\tilde{\mathcal{K}}}\coloneqq\bigl{\|}\bigl{(}\mathrm{e}^{\eta(\cdot)}v,\mathrm{e}^{\eta(\cdot)}Q,\mathrm{e}^{\eta(\cdot)}\ell,\mathrm{e}^{\eta(\cdot)}\omega,\mathrm{e}^{\eta(\cdot)}p\bigr{)}\bigr{\|}_{\tilde{\mathbb{E}}_{\infty}}.$
For $\varepsilon>0$, the ball with center zero and radius $\varepsilon$ in $\tilde{\mathcal{K}}$ is denoted by $\tilde{\mathcal{K}}_{\varepsilon}$, and we define $\Psi\colon\tilde{\mathcal{K}}_{\varepsilon}\to\tilde{\mathcal{K}}$ to be the solution map to (8.1). This means that given $(z,p)=(v,Q,\ell,\omega,p)\in\tilde{\mathcal{K}}_{\varepsilon}$, we denote by $\Psi(z,p)$ the solution to (8.1) with $G_{1}(z,p)$, $G_{2}(z)$, $G_{3}(z,p)$ and $G_{4}(z,p)$. By Theorem 6.1(b), the map $\Psi$ is well-defined if the terms on the right-hand side lie in $\mathrm{e}^{-\eta(\cdot)}\mathbb{F}_{\infty}\coloneqq\{z=(v,Q,\ell,\omega)\in\mathbb{F}_{\infty}:\mathrm{e}^{\eta(\cdot)}z\in\mathbb{F}_{\infty}\}$, where $\mathbb{F}_{\infty}$ is defined in (6.1), and $z_{0}=(v_{0},Q_{0},\ell_{0},\omega_{0})\in\mathrm{X}_{\gamma}$, with $\mathrm{X}_{\gamma}$ from (2.10). Throughout the section, we use $\|\cdot\|_{\infty}$ to denote the $\mathrm{L}^{\infty}(0,\infty)$-norm, while $\|\cdot\|_{\infty,\infty}$ represents the norm of $\mathrm{L}^{\infty}(0,\infty;\mathrm{L}^{\infty}(\mathcal{F}_{0}))$.
@@
[…]
Recall $X$, $Y$ as well as $g^{ij}$, $g_{ij}$ and $\Gamma_{jk}^{i}$ from Section 5, and let $(v,Q,\ell,\omega,p)\in\tilde{\mathcal{K}}_{\varepsilon}$. Then
@@
In the following, we take into account small enough $\varepsilon_{0}$ such that $\mathrm{J}_{X}$ is invertible for all $\varepsilon\in(0,\varepsilon_{0})$ by Section 8(a). The next lemma provides estimates of some of the terms appearing in (8.2). Here Section 8(a) follows as in [9, Proposition 7.4], while for the estimates in (b), we invoke Section 8(d) and proceed as in the proof of Section 7 for the estimate of $(v\cdot\nabla)Q$. Concerning Section 8(c), we make use of $(\mathrm{T}-\mathcal{T})(v,p,Q)=(\nabla v+(\nabla v)^{\top})(\operatorname{Id}_{3}-\mathbb{O})-\frac{2\xi}{3}(\operatorname{Id}_{3}-\mathbb{O}^{\top})(\Delta Q-Q)-\frac{2\xi}{3}\mathbb{O}^{\top}(\Delta Q-Q)(\operatorname{Id}_{3}-\mathbb{O})$, recall the mapping properties of $\mathcal{J}$ from Section 7(b) and employ Section 8(e).
[…]
Let $p>4$, and for $\varepsilon\in(0,\varepsilon_{0})$, consider $(v,Q,\ell,\omega,p)\in\tilde{\mathcal{K}}_{\varepsilon}$. Then there exists $C>0$ so that
@@
    $\|\mathrm{e}^{\eta(\cdot)}(\mathcal{L}_{1}-\Delta)v\|_{\mathbb{F}_{\infty}^{v}}$, $\|\mathrm{e}^{\eta(\cdot)}\mathcal{M}v\|_{\mathbb{F}_{\infty}^{v}}$, $\|\mathrm{e}^{\eta(\cdot)}\mathcal{N}(v)\|_{\mathbb{F}_{\infty}^{v}}$, $\|\mathrm{e}^{\eta(\cdot)}(\mathcal{G}-\nabla)p\|_{\mathbb{F}_{\infty}^{v}}\leq C\varepsilon^{2}$,
@@
    for $\mathcal{J}$ as defined in Section 7(b), it is valid that $\|\mathrm{e}^{\eta(\cdot)}\mathcal{J}((\mathrm{T}-\mathcal{T})(v,p,Q))\|_{p}\leq C\varepsilon^{2}$.
@@
[…]
Let $p>4$. Then there is $C>0$ so that for $\varepsilon\in(0,\varepsilon_{0})$ and $(z,p)$, $(z_{1},p_{1})$, $(z_{2},p_{2})\in\tilde{\mathcal{K}}_{\varepsilon}$, it holds that $\|\mathrm{e}^{\eta(\cdot)}G_{1}(z,p)\|_{\mathbb{F}_{\infty}^{v}}\leq C\varepsilon^{2}$ and $\|\mathrm{e}^{\eta(\cdot)}(G_{1}(z_{1},p_{1})-G_{1}(z_{2},p_{2}))\|_{\mathbb{F}_{\infty}^{v}}\leq C\varepsilon\cdot\|(z_{1},p_{1})-(z_{2},p_{2})\|_{\tilde{\mathcal{K}}}$.
@@
For the estimate of the linear term $\mathcal{B}_{\tau_{\mathrm{h},1}}$ from (5.8), we invoke the term $\operatorname{div}S_{\xi}(0)(\Delta-\operatorname{Id}_{3})Q$. Recalling that it is given by $-\nicefrac{{2\xi}}{{3}}\operatorname{div}(\Delta Q-Q)$, we expand the difference $\mathcal{B}_{\tau_{\mathrm{h},1}}(Q)-\operatorname{div}S_{\xi}(0)(\Delta-\operatorname{Id}_{3})Q$ as
@@
[…]
Assume $p>4$. There exists $C>0$ so that for $\varepsilon\in(0,\varepsilon_{0})$, $(z,p)$, $(z_{1},p_{1})$, $(z_{2},p_{2})\in\tilde{\mathcal{K}}_{\varepsilon}$, we have $\|\mathrm{e}^{\eta(\cdot)}G_{2}(z)\|_{\mathbb{F}_{\infty}^{Q}}\leq C\varepsilon^{2}$ and $\|\mathrm{e}^{\eta(\cdot)}(G_{2}(z_{1})-G_{2}(z_{2}))\|_{\mathbb{F}_{\infty}^{Q}}\leq C\varepsilon\cdot\|(z_{1},p_{1})-(z_{2},p_{2})\|_{\tilde{\mathcal{K}}}$.
@@
[…]
Let $p>4$. Then there is $C>0$ so that for all $\varepsilon\in(0,\varepsilon_{0})$, $(z,p)$, $(z_{1},p_{1})$, $(z_{2},p_{2})\in\tilde{\mathcal{K}}_{\varepsilon}$, it is valid that
$\|\mathrm{e}^{\eta(\cdot)}G_{3}(z,p)\|_{\mathbb{F}_{\infty}^{\ell}}$ $\leq C\varepsilon^{2}\enspace\text{and}\enspace\|\mathrm{e}^{\eta(\cdot)}(G_{3}(z_{1},p_{1})-G_{3}(z_{2},p_{2}))\|_{\mathbb{F}_{\infty}^{\ell}}\leq C\varepsilon\cdot\|(z_{1},p_{1})-(z_{2},p_{2})\|_{\tilde{\mathcal{K}}}\enspace\text{as well as}\enspace$
$\|\mathrm{e}^{\eta(\cdot)}G_{4}(z,p)\|_{\mathbb{F}_{\infty}^{\omega}}$ $\leq C\varepsilon^{2}\enspace\text{and}\enspace\|\mathrm{e}^{\eta(\cdot)}(G_{4}(z_{1},p_{1})-G_{4}(z_{2},p_{2}))\|_{\mathbb{F}_{\infty}^{\omega}}\leq C\varepsilon\cdot\|(z_{1},p_{1})-(z_{2},p_{2})\|_{\tilde{\mathcal{K}}}.$
@@
It readily follows that the terms $-m_{\mathcal{S}}\omega\times\ell$ as well as $\omega\times(J_{0}\omega)$ can be estimated by $C\varepsilon^{2}$. With regard to (8.2) and Section 8, it remains to discuss the term $-\int_{\partial\mathcal{S}_{0}}\mathbb{O}(t)^{\top}(\tau_{\mathrm{r}}(Q)+\sigma(Q))\mathbb{O}(t)n\,\mathrm{d}\Gamma$ in $G_{3}$ and the analogous term in $G_{4}$. As in the proof of Section 7, we focus on selected terms. The other terms can be treated in the same way. We have already discussed the linear part of $\tau$ in Section 8, and we observe that every remaining term has at least two factors of terms with $Q$, allowing for an estimate by $C\varepsilon^{2}$. Similarly as in the proof of Section 7, with the respective embeddings applied with $T=\infty$, and employing Section 8(e) for the boundedness of $\mathbb{O}$ and $\mathbb{O}^{\top}$, for small $\delta>0$, we get $\left\|\mathrm{e}^{\eta(\cdot)}\int_{\partial\mathcal{S}_{0}}(\mathbb{O}^{\top}\nabla Q\odot\nabla Q\mathbb{O})n\,\mathrm{d}\Gamma\right\|_{p}\leq C\cdot\|Q\|_{\mathrm{L}^{\infty}(0,\infty;\mathrm{H}^{\nicefrac{{3}}{{2}}+\delta}(\mathcal{F}_{0}))}\cdot\|\mathrm{e}^{\eta(\cdot)}Q\|_{\mathrm{L}^{p}(0,\infty;\mathrm{H}^{\nicefrac{{3}}{{2}}+\delta}(\mathcal{F}_{0}))}\leq C\varepsilon^{2}$, $\left\|\mathrm{e}^{\eta(\cdot)}\int_{\partial\mathcal{S}_{0}}(\mathbb{O}^{\top}Q\operatorname{tr}(Q\Delta Q)\mathbb{O})n\,\mathrm{d}\Gamma\right\|_{p}\leq C\varepsilon^{3}$ and $\left\|\mathrm{e}^{\eta(\cdot)}\int_{\partial\mathcal{S}_{0}}(\mathbb{O}^{\top}Q\operatorname{tr}(Q^{2})\operatorname{tr}(Q^{2})\mathbb{O})n\,\mathrm{d}\Gamma\right\|_{p}\leq C\varepsilon^{5}$. Hence, recalling $\mathcal{J}$ from Section 7(b), we deduce that $\|\mathrm{e}^{\eta(\cdot)}\mathcal{J}(-\mathbb{O}^{\top}(\tau_{\mathrm{r}}(Q)+\sigma(Q))\mathbb{O})\|_{p}\leq C\varepsilon^{2}$, so the first part of the assertion is implied. The Lipschitz estimate can be obtained likewise. ∎

### v2

@@ 7. Local strong well-posedness for large data
$\|\Phi_{T}^{R}(\tilde{z}_{1},\tilde{\pi}_{1})-\Phi_{T}^{R}(\tilde{z}_{2},\tilde{\pi}_{2})\|_{\tilde{\mathbb{E}}_{T}}\leq C_{\mathrm{MR}}L(R,T)\cdot\|(\tilde{z}_{1},\tilde{\pi}_{1})-(\tilde{z}_{2},\tilde{\pi}_{2})\|_{\tilde{\mathbb{E}}_{T}},\kern 5.0pt\text{for}\kern 5.0pt(\tilde{z}_{1},\tilde{\pi}_{1}),\,(\tilde{z}_{2},\tilde{\pi}_{2})\in\mathcal{K}_{T}^{R}.$
As $C_{\mathrm{MR}}>0$ is $T$-independent by virtue of the homogeneous initial values, we may choose $R>0$ and then also $T>0$ sufficiently small such that $C_{\mathrm{MR}}C(R,T)\leq\nicefrac{{R}}{{2}}$ and $C_{\mathrm{MR}}L(R,T)\leq\nicefrac{{1}}{{2}}$. In other words, the solution map $\Phi_{T}^{R}$ to the linearized problem (7.1) is a self-map and contraction, giving rise to a unique fixed point $(\widehat{z},\widehat{\pi})\in\prescript{}{0}{\mathbb{E}_{T}}\times\mathbb{E}_{T}^{\pi}$. An addition of the reference solution $(z^{*},\pi^{*})$ from Section 7 leads to the unique solution $(z,\pi)=(\widehat{z}+z^{*},\widehat{\pi}+\pi^{*})\in\tilde{\mathbb{E}}_{T}$ to (5.11)–(5.13). The proof is completed upon noting that $(z,\pi)\in\tilde{\mathbb{E}}_{T}$ means that $(v,Q,\ell,\omega,\pi)$ is in the regularity class as stated in Theorem 5.1.
@@ 8. Global strong well-posedness for small data
(8.1) $\left\{\begin{aligned} \partial_{t}{v}-\Delta{v}+\nabla\pi-\operatorname{div}S_{\xi}(0)(\Delta-\operatorname{Id}_{3})Q&=G_{1}(z,\pi),\kern 5.0pt\operatorname{div}{v}=0,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\mathcal{F}_{0},\\ {\partial_{t}{Q}}+\tilde{S}_{\xi}(0)\nabla v-(\Delta-\operatorname{Id}_{3}){Q}&=G_{2}(z),&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\mathcal{F}_{0},\\ m_{\mathcal{S}}({\ell})^{\prime}+\int_{\partial\mathcal{S}_{0}}\mathrm{T}(v,\pi,Q)n\,\mathrm{d}\Gamma&=G_{3}(z,\pi),&&\kern 5.0pt\text{in}\kern 5.0pt(0,T),\\ J_{0}({\omega})^{\prime}+\int_{\partial\mathcal{S}_{0}}y\times\mathrm{T}(v,\pi,Q)n\,\mathrm{d}\Gamma&=G_{4}(z,\pi),&&\kern 5.0pt\text{in}\kern 5.0pt(0,T),\\ {v}={\ell}+{\omega}\times y,\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\mathcal{S}_{0},\kern 5.0pt{v}&=0,\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\mathcal{O},\kern 5.0pt\partial_{n}Q=0,&&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\mathcal{F}_{0},\\ v(0)=v_{0},\kern 5.0ptQ(0)=Q_{0},\kern 5.0pt\ell(0)&=\ell_{0}\kern 5.0pt\text{and}\kern 5.0pt\omega(0)=\omega_{0}.\end{aligned}\right$.
@@
(8.2) $G_{1}(z,\pi)$ $\coloneqq(\mathcal{L}_{1}-\Delta)v-\mathcal{M}v-\mathcal{N}(v)-(\mathcal{G}-\nabla)\pi+\mathcal{B}_{\tau}(Q)+\mathcal{B}_{\sigma}(Q)-\operatorname{div}S_{\xi}(0)(\Delta-\operatorname{Id}_{3})Q,$
@@
$G_{3}(z,\pi)$ $\coloneqq-m_{\mathcal{S}}\omega\times\ell+\int_{\partial\mathcal{S}_{0}}\bigl((\mathrm{T}-\mathcal{T})(v,\pi,Q)-\mathbb{O}(t)^{\top}(\tau_{\mathrm{r}}(Q)+\sigma(Q))\mathbb{O}(t)\bigr)n\,\mathrm{d}\Gamma\kern 5.0pt\text{and}\kern 5.0pt$
$G_{4}(z,\pi)$ $\coloneqq\omega\times(J_{0}\omega)+\int_{\partial\mathcal{S}_{0}}y\times\bigl((\mathrm{T}-\mathcal{T})(v,\pi,Q)-\mathbb{O}(t)^{\top}(\tau_{\mathrm{r}}(Q)+\sigma(Q))\mathbb{O}(t)\bigr)n\,\mathrm{d}\Gamma.$
For $\eta\in(0,\eta_{0})$, where $\eta_{0}>0$ has been introduced in Theorem 6.1(b), and $\tilde{\mathbb{E}}_{\infty}$ as made precise in (6.3), we define $\tilde{\mathcal{K}}$ to be $\tilde{\mathbb{E}}_{\infty}$ with exponential weights, i. e., $(z,\pi)=(v,Q,\ell,\omega,\pi)\in\tilde{\mathcal{K}}$ if and only if
(8.3) $\mathrm{e}^{\eta(\cdot)}(v,Q,\ell,\omega,\pi)\in\tilde{\mathbb{E}}_{\infty},\kern 5.0pt\text{with}\kern 5.0pt\|(v,Q,\ell,\omega,\pi)\|_{\tilde{\mathcal{K}}}\coloneqq\bigl\|\bigl(\mathrm{e}^{\eta(\cdot)}v,\mathrm{e}^{\eta(\cdot)}Q,\mathrm{e}^{\eta(\cdot)}\ell,\mathrm{e}^{\eta(\cdot)}\omega,\mathrm{e}^{\eta(\cdot)}\pi\bigr)\bigr\|_{\tilde{\mathbb{E}}_{\infty}}$.
For $\varepsilon>0$, the ball with center zero and radius $\varepsilon$ in $\tilde{\mathcal{K}}$ is denoted by $\tilde{\mathcal{K}}_{\varepsilon}$, and we define $\Psi\colon\tilde{\mathcal{K}}_{\varepsilon}\to\tilde{\mathcal{K}}$ to be the solution map to (8.1). This means that given $(z,\pi)=(v,Q,\ell,\omega,\pi)\in\tilde{\mathcal{K}}_{\varepsilon}$, we denote by $\Psi(z,\pi)$ the solution to (8.1) with $G_{1}(z,\pi)$, $G_{2}(z)$, $G_{3}(z,\pi)$ and $G_{4}(z,\pi)$. By Theorem 6.1(b), the map $\Psi$ is well-defined if the terms on the right-hand side lie in $\mathrm{e}^{-\eta(\cdot)}\mathbb{F}_{\infty}\coloneqq\{z=(v,Q,\ell,\omega)\in\mathbb{F}_{\infty}:\mathrm{e}^{\eta(\cdot)}z\in\mathbb{F}_{\infty}\}$, where $\mathbb{F}_{\infty}$ is defined in (6.1), and $z_{0}=(v_{0},Q_{0},\ell_{0},\omega_{0})\in\mathrm{X}_{\gamma}$, with $\mathrm{X}_{\gamma}$ from (2.10). Throughout the section, we use $\|\cdot\|_{\infty}$ to denote the $\mathrm{L}^{\infty}(0,\infty)$-norm, while $\|\cdot\|_{\infty,\infty}$ represents the norm of $\mathrm{L}^{\infty}(0,\infty;\mathrm{L}^{\infty}(\mathcal{F}_{0}))$.
@@
Lemma 8.1.
Recall $X$, $Y$ as well as $g^{ij}$, $g_{ij}$ and $\Gamma_{jk}^{i}$ from Section 5, and let $(v,Q,\ell,\omega,\pi)\in\tilde{\mathcal{K}}_{\varepsilon}$. Then
@@
In the following, we take into account small enough $\varepsilon_{0}$ such that $\mathrm{J}_{X}$ is invertible for all $\varepsilon\in(0,\varepsilon_{0})$ by Section 8(a). The next lemma provides estimates of some of the terms appearing in (8.2). Here Section 8(a) follows as in [9, Proposition 7.4], while for the estimates in (b), we invoke Section 8(d) and proceed as in the proof of Section 7 for the estimate of $(v\cdot\nabla)Q$. Concerning Section 8(c), we make use of $(\mathrm{T}-\mathcal{T})(v,\pi,Q)=(\nabla v+(\nabla v)^{\top})(\operatorname{Id}_{3}-\mathbb{O})-\frac{2\xi}{3}(\operatorname{Id}_{3}-\mathbb{O}^{\top})(\Delta Q-Q)-\frac{2\xi}{3}\mathbb{O}^{\top}(\Delta Q-Q)(\operatorname{Id}_{3}-\mathbb{O})$, recall the mapping properties of $\mathcal{J}$ from Section 7(b) and employ Section 8(e).
Lemma 8.2.
Let $p>4$, and for $\varepsilon\in(0,\varepsilon_{0})$, consider $(v,Q,\ell,\omega,\pi)\in\tilde{\mathcal{K}}_{\varepsilon}$. Then there exists $C>0$ so that
@@
    $\|\mathrm{e}^{\eta(\cdot)}(\mathcal{L}_{1}-\Delta)v\|_{\mathbb{F}_{\infty}^{v}}$, $\|\mathrm{e}^{\eta(\cdot)}\mathcal{M}v\|_{\mathbb{F}_{\infty}^{v}}$, $\|\mathrm{e}^{\eta(\cdot)}\mathcal{N}(v)\|_{\mathbb{F}_{\infty}^{v}}$, $\|\mathrm{e}^{\eta(\cdot)}(\mathcal{G}-\nabla)\pi\|_{\mathbb{F}_{\infty}^{v}}\leq C\varepsilon^{2}$,
@@
    for $\mathcal{J}$ as defined in Section 7(b), it is valid that $\|\mathrm{e}^{\eta(\cdot)}\mathcal{J}((\mathrm{T}-\mathcal{T})(v,\pi,Q))\|_{p}\leq C\varepsilon^{2}$.
@@
Proposition 8.3.
Let $p>4$. Then there is $C>0$ so that for $\varepsilon\in(0,\varepsilon_{0})$ and $(z,\pi)$, $(z_{1},\pi_{1})$, $(z_{2},\pi_{2})\in\tilde{\mathcal{K}}_{\varepsilon}$, it holds that $\|\mathrm{e}^{\eta(\cdot)}G_{1}(z,\pi)\|_{\mathbb{F}_{\infty}^{v}}\leq C\varepsilon^{2}$ and $\|\mathrm{e}^{\eta(\cdot)}(G_{1}(z_{1},\pi_{1})-G_{1}(z_{2},\pi_{2}))\|_{\mathbb{F}_{\infty}^{v}}\leq C\varepsilon\cdot\|(z_{1},\pi_{1})-(z_{2},\pi_{2})\|_{\tilde{\mathcal{K}}}$.
@@
For the estimate of the linear term $\mathcal{B}_{\tau_{\mathrm{h},1}}$ from (5.8), we invoke the term $\operatorname{div}S_{\xi}(0)(\Delta-\operatorname{Id}_{3})Q$. Recalling that it is given by $\nicefrac{{2\xi}}{{3}}\operatorname{div}(\Delta Q-Q)$, we expand the difference $\mathcal{B}_{\tau_{\mathrm{h},1}}(Q)-\operatorname{div}S_{\xi}(0)(\Delta-\operatorname{Id}_{3})Q$ as
@@
Proposition 8.4.
Assume $p>4$. There exists $C>0$ so that for $\varepsilon\in(0,\varepsilon_{0})$, $(z,\pi)$, $(z_{1},\pi_{1})$, $(z_{2},\pi_{2})\in\tilde{\mathcal{K}}_{\varepsilon}$, we have $\|\mathrm{e}^{\eta(\cdot)}G_{2}(z)\|_{\mathbb{F}_{\infty}^{Q}}\leq C\varepsilon^{2}$ and $\|\mathrm{e}^{\eta(\cdot)}(G_{2}(z_{1})-G_{2}(z_{2}))\|_{\mathbb{F}_{\infty}^{Q}}\leq C\varepsilon\cdot\|(z_{1},\pi_{1})-(z_{2},\pi_{2})\|_{\tilde{\mathcal{K}}}$.
@@
Proposition 8.5.
Let $p>4$. Then there is $C>0$ so that for all $\varepsilon\in(0,\varepsilon_{0})$, $(z,\pi)$, $(z_{1},\pi_{1})$, $(z_{2},\pi_{2})\in\tilde{\mathcal{K}}_{\varepsilon}$, it is valid that
$\|\mathrm{e}^{\eta(\cdot)}G_{3}(z,\pi)\|_{\mathbb{F}_{\infty}^{\ell}}$ $\leq C\varepsilon^{2}\kern 5.0pt\text{and}\kern 5.0pt\|\mathrm{e}^{\eta(\cdot)}(G_{3}(z_{1},\pi_{1})-G_{3}(z_{2},\pi_{2}))\|_{\mathbb{F}_{\infty}^{\ell}}\leq C\varepsilon\cdot\|(z_{1},\pi_{1})-(z_{2},\pi_{2})\|_{\tilde{\mathcal{K}}}\kern 5.0pt\text{as well as}\kern 5.0pt$
$\|\mathrm{e}^{\eta(\cdot)}G_{4}(z,\pi)\|_{\mathbb{F}_{\infty}^{\omega}}$ $\leq C\varepsilon^{2}\kern 5.0pt\text{and}\kern 5.0pt\|\mathrm{e}^{\eta(\cdot)}(G_{4}(z_{1},\pi_{1})-G_{4}(z_{2},\pi_{2}))\|_{\mathbb{F}_{\infty}^{\omega}}\leq C\varepsilon\cdot\|(z_{1},\pi_{1})-(z_{2},\pi_{2})\|_{\tilde{\mathcal{K}}}$.
@@
It readily follows that the terms $m_{\mathcal{S}}\omega\times\ell$ as well as $\omega\times(J_{0}\omega)$ can be estimated by $C\varepsilon^{2}$. With regard to (8.2) and Section 8, it remains to discuss the term $\int_{\partial\mathcal{S}_{0}}\mathbb{O}(t)^{\top}(\tau_{\mathrm{r}}(Q)+\sigma(Q))\mathbb{O}(t)n\,\mathrm{d}\Gamma$ in $G_{3}$ and the analogous term in $G_{4}$. As in the proof of Section 7, we focus on selected terms. The other terms can be treated in the same way. We have already discussed the linear part of $\tau$ in Section 8, and we observe that every remaining term has at least two factors of terms with $Q$, allowing for an estimate by $C\varepsilon^{2}$. Similarly as in the proof of Section 7, with the respective embeddings applied with $T=\infty$, and employing Section 8(e) for the boundedness of $\mathbb{O}$ and $\mathbb{O}^{\top}$, for small $\delta>0$, we get $\left\|\mathrm{e}^{\eta(\cdot)}\int_{\partial\mathcal{S}_{0}}(\mathbb{O}^{\top}\nabla Q\odot\nabla Q\mathbb{O})n\,\mathrm{d}\Gamma\right\|_{p}\leq C\cdot\|Q\|_{\mathrm{L}^{\infty}(0,\infty;\mathrm{H}^{\nicefrac{{3}}{{2}}+\delta}(\mathcal{F}_{0}))}\cdot\|\mathrm{e}^{\eta(\cdot)}Q\|_{\mathrm{L}^{p}(0,\infty;\mathrm{H}^{\nicefrac{{3}}{{2}}+\delta}(\mathcal{F}_{0}))}\leq C\varepsilon^{2}$, $\left\|\mathrm{e}^{\eta(\cdot)}\int_{\partial\mathcal{S}_{0}}(\mathbb{O}^{\top}Q\operatorname{tr}(Q\Delta Q)\mathbb{O})n\,\mathrm{d}\Gamma\right\|_{p}\leq C\varepsilon^{3}$ and $\left\|\mathrm{e}^{\eta(\cdot)}\int_{\partial\mathcal{S}_{0}}(\mathbb{O}^{\top}Q\operatorname{tr}(Q^{2})\operatorname{tr}(Q^{2})\mathbb{O})n\,\mathrm{d}\Gamma\right\|_{p}\leq C\varepsilon^{5}$. Hence, recalling $\mathcal{J}$ from Section 7(b), we deduce that $\|\mathrm{e}^{\eta(\cdot)}\mathcal{J}(-\mathbb{O}^{\top}(\tau_{\mathrm{r}}(Q)+\sigma(Q))\mathbb{O})\|_{p}\leq C\varepsilon^{2}$, so the first part of the assertion is implied. The Lipschitz estimate can be obtained likewise. ∎